OpenStudy (anonymous):

x is a normally distributed random variable with a mean of 1, and p(-4

OpenStudy (zarkon):

p(-4<x<-2)=p(4<x<6)

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

if so , how you figure?

OpenStudy (zarkon):

it is not the final answer...it will lead you to the answer though

OpenStudy (anonymous):

how if It doesnt show me the standard deviation

OpenStudy (zarkon):

you don't need the standard deviation to answer this problem

OpenStudy (anonymous):

all I know is that the intervals |-4,-2| |4,6| are symmetric, but I cant figure the answer.

OpenStudy (zarkon):

do you agree that p(-4<x<-2)=p(4<x<6)?

OpenStudy (anonymous):

no because one is negative and the other is positive. right?

OpenStudy (zarkon):

well they are the same...using this fact we can get ... p(-4<x<-2)+p(4<x<6)=.2 p(-4<x<-2)+p(-4<x<-2)=.2 2*p(-4<x<-2)=.2 p(-4<x<-2)=.1

OpenStudy (anonymous):

I have to find that formula. That makes so much sense. thanks.

OpenStudy (anonymous):

is this the final answer?